OpenStudy (anonymous):

What is the missing expression? x^8 + 3x^5 = x^5(_____) 1.)x + 3 2.)x^2 + 3 3.)x^3 + 3

jimthompson5910 (jim_thompson5910):

First rewrite each term to be factored with the GCF as one of the factors. The GCF is \(\Large x^5\). Then factor out the GCF using the distributive property. \[\Large x^8 + 3x^5 = x^{5+3} + 3x^5\] \[\Large x^8 + 3x^5 = x^5 *x^3 + x^5 * 3\] \[\Large x^8 + 3x^5 = x^5 (x^3 + 3)\]